# FAST: Feature Aware Similarity Thresholding for Weak Unlearning in Black-Box Generative Models

### v1

@@ V-B4 Filtering Mechanism:
Accuracy Recall AUC Score FID Density Coverage Accuracy Recall AUC Score FID Density Coverage
Base Classifier 0.41$\pm$0.03 0.20$\pm$0.01 0.69$\pm$0.01 10.58$\pm$0.42 1$\pm$0.01 0.88$\pm$0.01 0.64$\pm$0.02 0.27$\pm$0.01 0.45$\pm$0.01 6.37$\pm$0.51 1.08$\pm$0.02 0.99$\pm$0.00
Base Classifier + Data Aug. 0.97$\pm$0.01 0.10$\pm$0.00 0.50$\pm$0.00 0.08$\pm$0.02 0.99$\pm$0.01 1$\pm$0.00 0.67$\pm$0.01 0.1$\pm$0.00 0.5$\pm$0.00 6.57$\pm$0.04 1.05$\pm$0.01 1$\pm$0.00
Imp-LS + MD 0.78$\pm$0.12 0.91$\pm$0.05 0.90$\pm$0.04 3.71$\pm$1.12 0.99$\pm$0.02 0.94$\pm$0.02 0.76$\pm$0.11 0.77$\pm$0.06 0.84$\pm$0.02 3.55$\pm$1.01 0.98$\pm$0.01 0.93$\pm$0.01
Imp-LS + MD + Latent Aug. 0.77$\pm$0.10 0.91$\pm$0.07 0.89$\pm$0.04 3.69$\pm$0.97 0.99$\pm$0.02 0.94$\pm$0.02 0.77$\pm$0.12 0.78$\pm$0.06 0.85$\pm$0.03 3.50$\pm$0.92 0.98$\pm$0.01 0.93$\pm$0.01
Imp-LS + Norm-SVM 0.78$\pm$0.08 0.93$\pm$0.06 0.92$\pm$0.02 2.54$\pm$1.00 0.99$\pm$0.01 0.96$\pm$0.02 0.77$\pm$0.07 0.82$\pm$0.04 0.86$\pm$0.02 3.81$\pm$0..93 0.95$\pm$0.01 0.93$\pm$0.01
Imp-LS + Norm-SVM + Latent Aug. 0.81$\pm$0.06 0.93$\pm$0.02 0.93$\pm$0.05 2.31$\pm$0.72 0.99$\pm$0.01 0.97$\pm$0.01 0.76$\pm$0.05 0.77$\pm$0.03 0.84$\pm$0.05 3.35$\pm$0.65 0.99$\pm$0.15 0.94$\pm$0.01
Inv-LS + MD 0.71$\pm$0.02 0.89$\pm$0.01 0.89$\pm$0.01 9.91$\pm$1.14 0.96$\pm$0.01 0.85$\pm$0.01 0.78$\pm$0.06 0.86$\pm$0.03 0.88$\pm$0.02 9.06$\pm$0.72 0.94$\pm$0.01 0.83$\pm$0.02
Inv-LS + MD + Latent Aug. 0.72$\pm$0.02 0.89$\pm$0.01 0.89$\pm$0.01 9.86$\pm$1.04 0.96$\pm$0.01 0.85$\pm$0.01 0.79$\pm$0.06 0.86$\pm$0.03 0.89$\pm$0.02 9.05$\pm$0.70 0.94$\pm$0.01 0.83$\pm$0.02
Inv-LS + Norm-SVM 0.80$\pm$0.01 0.94$\pm$0.02 0.93$\pm$0.03 3.61$\pm$0.65 1.01$\pm$0.02 0.96$\pm$0.01 0.76$\pm$0.05 0.74$\pm$0.01 0.85$\pm$0.01 5.15$\pm$0.81 0.93$\pm$0.02 0.90$\pm$0.01
Inv-LS + Norm-SVM + Latent Aug. 0.80$\pm$0.02 0.92$\pm$0.02 0.93$\pm$0.02 3.83$\pm$0.70 1.00$\pm$0.01 0.94$\pm$0.01 0.78$\pm$0.05 0.78$\pm$0.01 0.86$\pm$0.01 5.48$\pm$0.75 0.94$\pm$0.02 0.90$\pm$0.01
V-C Baseline and Evaluation Metrics
Baselines: To the best of our knowledge, there exists no established baseline method for effectively filtering out undesired output samples generated by a black-box GAN. In response to this gap in the literature, our approach introduces a fundamental yet essential baseline strategy to address this challenge. Our methodology entails the development and deployment of a binary classifier, designed to leverage user feedback consisting of positive and negative samples. This classifier is tailored to identify and filter out undesired output samples, serving as a foundational component in our approach.
Evaluation Metrics: To gauge the effectiveness of our proposed techniques and the baseline methods, we utilize three fundamental evaluation metrics:
@@ V-C Baseline and Evaluation Metrics
    AUC-Score: This metric serves as a quantitative measure to assess the effectiveness of the filtering method. It is important to note that the highest achievable value for this metric is 1, indicating optimal performance.
@@
    Recall: In the context of this task, the recall value is a crucial metric as it measures the effectiveness of filtering out negative images generated by the model. It quantifies the proportion of negative images successfully filtered by the filtering method. An optimal recall value, represented by 1, indicates that all negative images have been correctly filtered out.
@@
    Fréchet Inception Distance (FID), Density and Coverage: While the earlier mentioned metrics quantify the effectiveness of filtering, they do not assess the quality of the filtered output distribution in comparison to a distribution obtained through human-level filtering. To replicate human-level filtering, we utilize a pre-trained classifier trained on labeled training data. To address this aspect, we calculate the Fréchet Inception Distance (FID) [46] between the generated samples and a classifier trained on labeled data. A lower FID value indicates a higher accuracy of the filtering mechanism, signifying a closer match to the true distribution. Additionally, we employ Density and Coverage [47] metrics to evaluate how the output of our proposed filtering method deviates from the distribution obtained with a pre-trained (human-level) classifier. The optimal value for Coverage is 1, and higher values of Density and Coverage signify better filtering performance.
Please note that the original dataset is unavailable during the filtering process. Consequently, the use of the original dataset is solely for evaluation purposes.
V-D Results
@@
Table I and Table II provide evaluations of different filtering methods on the MNIST and CelebA-HQ datasets, respectively.We observe that our proposed methods give better Recall and AUC values for both datasets while comparable to superior performance for other metrics.
@@
The analysis focuses on two specific class features, Class-5 and Class-8 for MNIST, and subtle features of Bangs and Hats for CelebA-HQ, each with only 20 positive and 20 negative user feedback samples. For both MNIST and CelebA-HQ, we observe that the Base Classifier, trained on the user feedback data, exhibits moderate performance in both classes. However, incorporating data augmentation significantly improves accuracy, with notable enhancements in the AUC score. Methods employing Implicit-Latent Space (Imp-LS) with Mean Difference (MD) or Normal SVM (Norm-SVM) feature representation-based filtering mechanism(Imp-LS+MD/Norm-SVM) demonstrate better performance than the Base Classifier in terms of Recall and AUC scores for both the datasets. The Imp-LS+MD and Imp-LS+MD+Feature-Augmentation methods demonstrate superior recall performance, surpassing the Base Classifier by approximately 3.5 times for MNIST Class-5 and 2 times for MNIST Class-8. Additionally, these methods achieve remarkable recall scores, outperforming the Base Classifier by almost 4.1 times for CelebA-HQ Bangs and 1.7 times for Hats features, respectively. The combination of Inversion-Latent Space (Inv-LS) with Normal SVM (Norm-SVM) feature representation-based mechanism (Inv-LS+Norm-SVM) yields superior AUC scores, demonstrating improvements of 66.4% and 35.8% over the base classifier for MNIST Class-5 and Class-8, respectively. In terms of FID, Density, and Coverage we observe that our proposed family of methods gives superior to comparative performance with respect to the Base Classifier.
These findings suggest the efficacy of incorporating model distillation and latent augmentation techniques, especially within an invertible framework, to enhance the performance of classifiers on classes with limited user feedback samples.
V-E Ablation Study
V-E1 Effect of user-feedback size
@@
Table III presents an evaluation of recall and AUC scores, examining the efficacy of diverse methods on two distinct features—MNIST Class-5 and CelebA-HQ Bangs—across varying numbers of positive and negative feedback samples.
@@
TABLE III: Evaluation of Recall and AUC scores with varying no of pos and negative feedback
No of Pos. and Neg. Samples Methods MNIST Class-5 CelebA-HQ - Bangs
@@
V-E2 Evaluation under no positive feedback
In our endeavor to enhance the adaptability of our approach, we extend its functionality to accommodate scenarios where users provide exclusively negative samples without any positive instances. In such cases, we employ a technique referred to as positive mining to infer positive samples. Specifically, we select positive samples based on the projection similarity of latent vectors with the negative latents provided by the user.
Table IV offers a comprehensive evaluation summary under the constraint of 20 negative and 0 positive feedback instances for MNIST Class-5 and CelebA Bangs datasets. Leveraging positive mining, our methods, Imp-LS + MD and Imp-LS + MD + Latent Augmentation, achieve noteworthy recall values of 0.72 and 0.73 for MNIST Class-5, and 0.98 and 0.99 for CelebA Bangs, respectively. Remarkably, these recall values surpass the performance achieved when both positive and negative feedback samples are available.
TABLE IV: Evaluation of Recall and AUC scores with only 20 negative and 0 positive feedback
Methods MNIST Class-5 CelebA - bangs
@@ V-E2 Evaluation under no positive feedback
VI Conclusion
In this research, we have explored algorithms designed to prevent the generation of samples that exhibit undesired features from a black-box pre-trained Generative Adversarial Network (GAN). Recognizing the inadequacy of current unlearning methods, primarily due to the unavailability of the pre-trained model’s parameters, we propose a filtering methodology as an alternative solution. Within the theoretical framework of this study, we establish that filtering can be seen as a manifestation of weak unlearning. It is noteworthy that our methodology operates effectively even when the original training dataset and model parameters remain concealed from the user, allowing it to extend to zero-shot settings. This work represents a pioneering endeavor that sheds light on the intricate relationship between filtering and unlearning, offering a fresh perspective on the challenges posed by black-box generative models. However, it’s essential to acknowledge certain limitations that our methodology may not address comprehensively, particularly in scenarios involving the simultaneous filtering of correlated features alongside undesired features. The intricate interplay of semantic features within images can occasionally lead to unintended consequences on other desirable attributes, thus necessitating further investigation in this direction.

### v2

@@ 5.2.2 Pre-trained Classifier
Recall AUC FID Density Coverage Recall AUC FID Density Coverage
Base Classifier 0.20$\pm$0.01 0.69$\pm$0.01 10.58$\pm$0.42 1$\pm$0.01 0.88$\pm$0.01 0.27$\pm$0.01 0.45$\pm$0.01 6.37$\pm$0.51 1.08$\pm$0.02 0.99$\pm$0.00
Base Classifier+Data Aug. 0.10$\pm$0.00 0.50$\pm$0.00 0.08$\pm$0.02 0.99$\pm$0.01 1$\pm$0.00 0.1$\pm$0.00 0.5$\pm$0.00 6.57$\pm$0.04 1.05$\pm$0.01 1$\pm$0.00
Imp-LS+MD 0.91$\pm$0.05 0.90$\pm$0.04 3.71$\pm$1.12 0.99$\pm$0.02 0.94$\pm$0.02 0.77$\pm$0.06 0.84$\pm$0.02 3.55$\pm$1.01 0.98$\pm$0.01 0.93$\pm$0.01
Imp-LS+MD+Latent Aug. 0.91$\pm$0.07 0.89$\pm$0.04 3.69$\pm$0.97 0.99$\pm$0.02 0.94$\pm$0.02 0.78$\pm$0.06 0.85$\pm$0.03 3.50$\pm$0.92 0.98$\pm$0.01 0.93$\pm$0.01
Imp-LS+Norm-SVM 0.93$\pm$0.06 0.92$\pm$0.02 2.54$\pm$1.00 0.99$\pm$0.01 0.96$\pm$0.02 0.82$\pm$0.04 0.86$\pm$0.02 3.81$\pm$0..93 0.95$\pm$0.01 0.93$\pm$0.01
Imp-LS+Norm-SVM+Latent Aug. 0.93$\pm$0.02 0.93$\pm$0.05 2.31$\pm$0.72 0.99$\pm$0.01 0.97$\pm$0.01 0.77$\pm$0.03 0.84$\pm$0.05 3.35$\pm$0.65 0.99$\pm$0.15 0.94$\pm$0.01
Inv-LS+MD 0.89$\pm$0.01 0.89$\pm$0.01 9.91$\pm$1.14 0.96$\pm$0.01 0.85$\pm$0.01 0.86$\pm$0.03 0.88$\pm$0.02 9.06$\pm$0.72 0.94$\pm$0.01 0.83$\pm$0.02
Inv-LS+MD+Latent Aug. 0.89$\pm$0.01 0.89$\pm$0.01 9.86$\pm$1.04 0.96$\pm$0.01 0.85$\pm$0.01 0.86$\pm$0.03 0.89$\pm$0.02 9.05$\pm$0.70 0.94$\pm$0.01 0.83$\pm$0.02
Inv-LS+Norm-SVM 0.94$\pm$0.02 0.93$\pm$0.03 3.61$\pm$0.65 1.01$\pm$0.02 0.96$\pm$0.01 0.74$\pm$0.01 0.85$\pm$0.01 5.15$\pm$0.81 0.93$\pm$0.02 0.90$\pm$0.01
Inv-LS+Norm-SVM+Latent Aug. 0.92$\pm$0.02 0.93$\pm$0.02 3.83$\pm$0.70 1.00$\pm$0.01 0.94$\pm$0.01 0.78$\pm$0.01 0.86$\pm$0.01 5.48$\pm$0.75 0.94$\pm$0.02 0.90$\pm$0.01
5.3 Baseline and Evaluation Metrics
5.3.1 Baseline
To the best of our knowledge, there exists no established baseline method for effectively filtering out undesired output samples generated by a black-box GAN. So to evaluate the effectiveness of our method, we implement a fundamental yet essential baseline strategy of using a binary classifier designed to leverage user feedback consisting of positive and negative samples. This classifier termed as Base Classifier is tailored to identify and filter out undesired output samples, serving as a foundational baseline in our approach. To implement this Base Classifier we have used ResNext50 model [44] pre-trained on Imagenet as the base feature extractor on top of which 3-layer MLP trained with binary cross entropy loss, fine-tuned for both MNIST and CelebA-HQ datasets. Adamax optimizer is used for training, running for 200 epochs with a learning rate of 0.07. Given the limited user feedback ($\leq 100$ samples), we introduce data augmentation techniques including rotation, blur, perspective, and auto-augment to boost the performance of this base classifier. Each image is augmented into 64 variations during training.
5.3.2 Evaluation Metrics
To gauge the effectiveness of our proposed techniques and the baseline methods, we utilize three fundamental evaluation metrics:
@@ 5.3.2 Evaluation Metrics
    Recall: In this task, recall is a crucial metric as it measures the effectiveness of filtering out negative images generated by the model. It quantifies the proportion of negative images successfully filtered. An optimal recall value of 1 indicates that all negative images have been correctly filtered out.
@@
    AUC: The AUC score, derived from varying the threshold $T_{th}$ and calculating the area under the curve of the True Positive rate ($P_{D}$) vs. the False Positive rate ($P_{F}$), quantitatively assesses the filtering method’s effectiveness. Thus AUC captures the effect of varying threshold. A maximum AUC score of 1 indicates optimal performance.
@@
    Fréchet Inception Distance (FID), Density and Coverage: While the earlier mentioned metrics quantify the effectiveness of filtering, they do not assess the quality of the filtered output distribution in comparison to a distribution obtained through human-level filtering. To replicate human-level filtering, we utilize a pre-trained classifier trained on labeled training data [^5] Please note that the original dataset is unavailable during the FAST filtering process. Consequently, the use of the original dataset is solely for evaluation purposes with the Pre-trained Classifier.. To address this aspect, we calculate the Fréchet Inception Distance (FID) [45] between the generated samples after the FAST method and a pre-trained classifier trained on labeled data. A lower FID value signifies a closer match to the pre-trained classifier’s output distribution. Additionally, we employ Density and Coverage [46] metrics to evaluate how the output of our proposed filtering method deviates from the distribution obtained with a pre-trained (human-level) classifier. The optimal value for Coverage is 1, and higher values of Density and Coverage signify better filtering performance.
5.4 Results
@@
Table-1 and Table-2 provide evaluations of different filtering methods on the MNIST and CelebA-HQ datasets, respectively. We observe that our proposed methods give better Recall and AUC values for both datasets while comparable to superior performance for other metrics.
@@
The analysis focuses on two specific class features, class-5 and class-8 for MNIST, and subtle features of bangs and hats for CelebA-HQ, each with only 20 positive and 20 negative user feedback samples. For both MNIST and CelebA-HQ, we observe that the Base Classifier, trained on the user feedback data, exhibits moderate performance in both classes. However, incorporating data augmentation significantly improves accuracy, with notable enhancements in the AUC score. Methods employing Implicit-Latent Space (Imp-LS) with Mean Difference (MD) or Normal SVM (Norm-SVM) feature representation-based filtering mechanism (Imp-LS+MD/Norm-SVM) demonstrate better performance than the Base Classifier in terms of Recall and AUC scores for both the datasets. The Imp-LS+MD and Imp-LS+MD+Latent-Augmentation methods demonstrate superior recall performance, surpassing the Base Classifier by approximately 3.5 times for MNIST class-5 and 2 times for MNIST class-8. Additionally, these methods achieve remarkable recall scores, outperforming the Base Classifier by almost 4.1 times for CelebA-HQ bangs and 1.7 times for hats features, respectively. The combination of Inversion-Latent Space (Inv-LS) with Normal SVM (Inv-LS+Norm-SVM) feature representation-based mechanism yields superior AUC scores, demonstrating improvements of 66.4% and 35.8% over the Base Classifier for MNIST class-5 and class-8, respectively. In terms of FID, Density, and Coverage we observe that our proposed family of methods gives superior to comparative performance with respect to the Base Classifier.
These findings suggest the efficacy of the FAST family of methods. Although no single method achieves superior results across all metrics, in most cases, Inv-LS+Norm-SVM and Inv-LS+Norm-SVM+Latent Aug. perform significantly better than other methods for both Recall and AUC scores. This makes them the most desirable methods among the proposed ones.
5.5 Ablation Study
5.5.1 Effect of user-feedback size
@@
Table-3 presents an evaluation of recall and AUC scores, examining the efficacy of diverse methods on two distinct features—MNIST class-5 and CelebA-HQ bangs—across varying numbers of positive and negative feedback samples.
@@
Table 3: Evaluation of Recall and AUC scores with varying no of pos and negative feedback
No. of (Pos. + Neg.) Samples Methods MNIST Class-5 CelebA-HQ - Bangs
@@
5.5.2 Evaluation under no positive feedback
In our endeavor to enhance the adaptability of our approach, we extend its functionality to accommodate scenarios where users provide exclusively negative samples without any positive instances. In such cases, we employ a technique referred to as positive mining to infer positive samples. Specifically, given the user identified negative samples $\mathcal{S}_{n}=\{y^{n}_{i}\}_{i=1}^{s}$, their corresponding negative latents $\mathcal{Z}_{n}=\{z^{n}_{i}\}_{i=1}^{s}$ are computed. The negative similarity scores set $SS(z^{n}_{i},\bar{z}^{n})=\{sim(z^{n}_{i},\bar{z}^{n}):z^{n}_{i}\in\mathcal{S}_{n},\bar{z}^{n}=\frac{1}{s}\sum^{s}_{i=1}z^{n}_{i}\}$ is defined. Now given a newly sampled latent $z\sim P_{\mathcal{Z}}$, it is a positive latent if $sim(z,\bar{z}^{n})\ll\min_{i}SS(z^{n}_{i},\bar{z}^{n})$. The results of positive mining are given in Fig. 3 for the MNIST dataset, using only 20 negative samples from Class-5 and Class-8. These results were generated by a pre-trained GAN after obtaining positive noises through the positive mining approach. It is evident that positive mining not only selects noise samples from other classes but also maintains diversity across these classes, demonstrating its effectiveness even when user feedback is limited to negative samples.
(a) pos. mining Class-5
(b) pos. mining Class-8
Figure 3: Results of positive mining given only 20 negative samples by the user for MNIST dataset. Samples are generated by the pre-trained GAN after obtaining the positive samples through positive mining.
Table-4 offers a comprehensive evaluation summary under the constraint of 20 negative and 0 positive feedback instances for MNIST class-5 and CelebA-HQ bangs datasets. Leveraging positive mining, our methods, Imp-LS+MD and Imp-LS+MD+Latent Augmentation, achieve noteworthy recall values of 0.72 and 0.73 for MNIST class-5, and 0.98 and 0.99 for CelebA-HQ bangs, respectively. Remarkably, these recall values surpass the performance achieved when both positive and negative feedback samples are available.
Table 4: Evaluation of Recall and AUC scores with only 20 negative and 0 positive feedback
Methods MNIST Class-5 CelebA-HQ - Bangs
@@ 5.5.2 Evaluation under no positive feedback
6 Conclusion, Limitations and Future Work
To deploy deep generative models safely and responsibly, it is crucial to eliminate outputs with undesired features. In this research, we explored algorithms to prevent generating such samples from a black-box pre-trained Generative Adversarial Network (GAN). Recognizing that current unlearning methods are inadequate due to the lack of access to the pre-trained model’s parameters, we propose a filtering methodology as an alternative solution. In our theoretical framework, we establish that filtering can be viewed as a form of weak unlearning. Thus to effectively filter undesired output, we propose an algorithm named Feature Aware Similarity Thresholding (FAST), which suppresses unwanted outputs by systematically encoding the representation of undesired features in the latent space. Notably, our methodology works effectively even when the original training dataset and model parameters are hidden from the user, making it applicable to zero-shot settings. This work is a pioneering effort that highlights the intricate relationship between filtering and unlearning, offering a new perspective on the challenges posed by black-box generative models.
Limitations: The primary inspiration for our method is the understanding that latent space efficiently encodes semantic features. Therefore, our goal is to represent the undesired feature within the latent space, which necessitates certain useful properties such as disentanglement between different features. Disentanglement ensures that distinct features are represented independently, allowing for precise filtering of undesired attributes. However, in many cases, the latent space may be entangled due to the intricate interplay of semantic features within images. This entanglement can lead to the representation of multiple correlated features in a combined manner, making it challenging to isolate and filter out only the undesired features. As a result, our methodology might inadvertently filter out correlated features along with the undesired features in such scenarios.
Future works: Acknowledging the limitations, concerns, and further applicability of our method, we have identified key aspects to address in future work: (i) Learning disentangled latent space: We plan to explore the use of inversion mechanisms to learn a more disentangled latent space, which would reduce the effect of feature entanglement. This approach aims to improve the precision of filtering undesired features without inadvertently affecting correlated features. (ii) Application of FAST to black-box diffusion models: We aim to extend the FAST algorithm to black-box diffusion models. We hypothesize that further modifications to the FAST mechanism will be required for diffusion models because these models generate samples through multiple Langevian backward steps. This process suggests that the initial latent space may not encode semantic features effectively, necessitating additional investigation and adaptation of the FAST algorithm to ensure its applicability and effectiveness in this context.